Please write neatlyLet f(x) be a Cl(R) function and u(t, x) be a solution to the inviscid Burgers IVP Ut + uuz = 0, 4(0,x) = f(x). (a) Use the
Answer #1

ANSWER Given that oco,x) = fre) langranges equation at = dd - du 3) 0 = constant =4, 0 and de = dt . a-cit = 2 - 0 from (1) a= -ofx-ut) itt plcx-ut) = -f(x-ut) f(X-ut) itt f(x-ut) 5-f1&Jfcę) Here $ = x-ut. H++ plie) b) XER, such that flexo) 20, a